OpenStudy (saaarrrahhhh):

Adele is 5 years older than Timothy. In three years, Timothy will be 2/3 of Adele's age. What is Adele's current age?

OpenStudy (amorfide):

Timothy will be of Adele’s age. What is Adele’s current age? did you forget to put something before of?

OpenStudy (saaarrrahhhh):

Yes thanks I fixed it! haha

OpenStudy (bibby):

a=t+5 t+3=a+3

OpenStudy (anonymous):

Here is a good way to solve this type of equation: First, find a common factor between the two expressions: Then, we know that the equation is using multiplication as it says 2/3 of Adele's age, so we multiply our finished values: